OpenStudy (anonymous):

What is 10^6 in standard form? A. 0.0000001 B. 0.000001 C. 1,000,000 D. 10,000,000

OpenStudy (anonymous):

C

jigglypuff314 (jigglypuff314):

10^3 = 1 with 3 zeros = 1000 10^7 = 1 with 7 zeros = 10000000 so 10^6 = ?

OpenStudy (kc_kennylau):

10^6=10*10*10*10*10*10=?

OpenStudy (anonymous):

thanks a loooot!!! :)

jigglypuff314 (jigglypuff314):

glad we could help :)